# When Is Diversity Rewarded in Cooperative Multi-Agent Learning?

### v1

@@ Appendix A Computational Resources Used
Simply verifying our results using the available code will take considerably less compute hours (around a day).
@@ Appendix B Code and Data Availability
We attach the code in the supplementary materials. The code contains instructions on how to reproduce the experiments in the paper and dedicated YAML files containing the hyperparameters for each experiment presented. The YAML files are structured according to the HYDRA [45] framework which allows smooth reproduction as well as systematic and standardized configuration. We further attach all scripts to reproduce the plots in the paper from the experiment results.
@@ Appendix C Implementation Details
For all experiments, we use the MAPPO MARL algorithm [44]. Environments are implemented in the multi-agent environment simulator VMAS [46], and trained using TorchRL [47]. Both the actor and critic are two-layer MLPs with 256 neurons per layer and Tanh activation. Further details, such as hyperparameter choices, are available in the attached code and YAML configuration files.
Appendix D Colonel Blotto & Level-Based Foraging
@@
D.1 Team Colonel Blotto (fixed adversary)
The Colonel Blotto game is a well-known allocation game studied in both game theory and MARL [7, 8]. It is used to model election strategies and other resource-based competitions. In the team variant with fixed adversary, $N$ friendly colonels (agents) each distribute a (fixed and equal) budget of troops $r_{ij}\!\geq 0,\sum_{j=1}^{M}r_{ij}=1$ across $M$ battlefields $j\in\{1,\dots,M\}$ (our tasks). A fixed adversary selects a stochastic opposing allocation strategy, i.e. a distribution $\pi_{\mathrm{adv}}$ over vectors $a=(a_{1},\dots,a_{M})$ which is fixed throughout training and evaluation. Let $s_{j}=\sum_{i=1}^{N}r_{ij}$ denote the team force committed to battlefield $j$. Our agents win against the adversary if the troops they allocate to a given field surpass the troops allocated by the adversary. The expected value secured on battlefield $j$ is therefore
@@ D.1 Team Colonel Blotto (fixed adversary)
D.2 Level-Based Foraging
The well-known level-based foraging (LBF) benchmark, based on the knapsack problem [48], is a deceptively challenging, embodied MARL environment, where $N$ agents are placed on a grid with $M$ food items, and are tasked with collecting them. Each item $j$ has an integer level $L_{j}$ that must be met or exceeded by the combined skills of the agents standing on that cell before it can be collected [9]. Let agent $i$’s skill be $e_{i}$. At a given step the binary variable
@@ D.2 Level-Based Foraging
Appendix E Sum-Form Aggregators
@@
Definition E.1 (Sum-Form Aggregator).
@@
Lemma E.1 (Schur Properties of Sum-Form Aggregators [40]).
@@
Corollary E.1 (Convex-Concave Positive Heterogeneity Gain).
@@
First, by Lemma E.1, since $g$ is strictly convex, the (identical) task-level aggregators $T_{j}(\mathbf{x})=\sum_{i=1}^{N}g(x_{i})$ are strictly Schur-convex, satisfying condition (i) of Theorem 3.1.
@@
Hence, the conditions of Thm. 3.1 apply. To establish $\Delta R>0$, it remains to verify that the optimal allocation is non-trivial: it distributes effort across at least two tasks. In any admissible homogeneous solution, each of the $N$ agents chooses the same effort-distribution $(c_{1},\dots,c_{M})$ on tasks, with $\sum_{j}c_{j}=1$. Then task $j$’s reward is $T_{j}=N\,g(c_{j})$, so $R(\mathbf{A})\;=\;\sum_{j=1}^{M}h\!\bigl{(}N\,g(c_{j})\bigr{)}.$ The trivial, all-agent single-task allocation uses $(c_{j}=1,c_{k\neq j}=0)$. Its reward is therefore $R_{\text{corner}}=h\!\bigl{(}N\,g(1)\bigr{)}+\sum_{k\neq j}h\bigl{(}N\,g(0)\bigr{)}=h\!\bigl{(}N\,g(1)\bigr{)}$ since $g(0)=0$ and $h(0)=0$.
@@
Appendix F Formal Analysis
F.1 Proof of Thm. 3.1
@@ Appendix F Formal Analysis
Then each column $j$ of $\mathbf{A}_{\mathrm{hom}}$ is the uniform vector $\mathbf{u}_{j}\;=\;(c_{j},c_{j},\dots,c_{j})^{\top}\;\in\mathbb{R}^{N}.$ Hence the task-level reward is $T_{j}(\mathbf{u}_{j})$, and the overall reward is
@@
The total effort allocated to a task $j$ can be expressed as $\lfloor Nc_{j}\rfloor+f_{k}$, where $10\leq f_{j}<$ 1. First, we assign $\lfloor Nc_{j}\rfloor$ agents to allocate effort $1$ to task $j$, for every task $j$. These agents are all distinct. This leaves us with $\sum_{j}f_{j}~{}=~{}N~{}-~{}\sum_{j}\lfloor Nc_{j}\rfloor$ agents that have not allocated any effort yet. Let $i$ be the first of those agents. We have agent $i$ allocate $f_{1}$ effort to task $1$, $f_{2}$ effort to task $2$, and so on, until we arrive at a task $k$ such that $f_{1}+\ldots+f_{k}=1+s$, for some $s>0$. We have $i$ allocate $f_{k}-s$ to this task $k$. Then, we move to agent $i+1$, and allocate the remaining fractional efforts in the same manner (and in particular, allocating $s$ effort to task $k$), until agent $i+1$ overflows. Then we move to agent $i+2$, and so on. This ensures that we have allocated $N$ effort in total across the agents, and that every agent’s effort allocation sums exactly to $1$, so is feasible.
@@
which is clearly homogeneous (each row is the same), and respects each column sum $s_{j}$. Since $\sum_{j}s_{j}=N$, each row sums to $1$, hence the allocation is feasible. By Schur-concavity of $T_{j}$, for each column $j$ we have
@@
unless $\mathbf{a}_{j}$ is $\mathbf{u}_{j}$. In other words, any deviation from the uniform vector with the same sum $\sum_{i=1}^{N}a_{ji}=s_{j}$ will not increase $T_{j}(\mathbf{a}_{j})$ under Schur-concavity. Hence for each column $j$ of $\mathbf{A}$, $T_{j}(\mathbf{a}_{j})\;\leq\;T_{j}(\mathbf{u}_{j}),$. Since $U$ is non-decreasing in each coordinate,
@@
[…]
@@
[…]
@@
Appendix G Deriving the $\{\min,\textnormal{mean},\max\}$ heterogeneity gains in the Fig. 2 table
@@
$U=$ mean Inner: $T_{j}=\min_{i}r_{ij}=1/M.$ Outer: simple avg $\tfrac{1}{M}\sum_{j}T_{j}.$ Since $\sum_{j}T_{j}=M\cdot(1/M)=1\implies R=1/M.$ Both $R_{het}$, $R_{hom}$ same $\implies\Delta R=0.$ Gap: 0. Inner: $T_{j}=\tfrac{1}{N}\sum_{i}r_{ij}.$ Then $\sum_{j}T_{j}=1.$ Outer: avg $\!=\!\tfrac{1}{M}\sum_{j}T_{j}.$ Hence $R=\tfrac{1}{M}\cdot 1=\tfrac{1}{M}.$ Same for $R_{het}$, $R_{hom}$. Gap: 0. Inner: $T_{j}=\max_{i}r_{ij}.$ Outer: avg $\!=\!\tfrac{1}{M}\sum_{j}T_{j}.$ $R_{het}$: sum $=M\implies R=1.$ $R_{hom}$: sum $=1\implies R=1/M.$ Gap: $1-\tfrac{1}{M}=\tfrac{M-1}{M}.$
@@
$\min$ Inner: $T_{j}\!\to\!\begin{cases}1,&\text{if all agents pick $j$},\\ 0,&\text{otherwise}.\end{cases}$ Outer: $\min_{j}T_{j}$. To get $R>0$, must have $T_{j}>0$ for every $j$ (i.e. all agents pick all tasks, impossible). Hence $R_{\mathrm{het}}=R_{\mathrm{hom}}=0$ typically, $\Delta R=0.$ Inner: $T_{j}=\frac{|\mathcal{I}_{j}|}{N}$ Outer: $\min_{j}T_{j}$. $R_{het}=\lfloor N/M\rfloor/N$. $R_{hom}=0$. $\Delta R=\lfloor N/M\rfloor/N$. Inner: $T_{j}\!\to\!\begin{cases}1,&\text{if at least 1 agent picks $j$},\\ 0,&\text{if no agent picks $j$}.\end{cases}$ Outer: $\min_{j}T_{j}.$ Heterogeneous can choose $s$ distinct tasks. If want $\min_{j}=1$, must pick all $M$ tasks. That requires $N\geq M$. Then $R=1.$ Homogeneous covers only 1 task $\implies\min_{j}=0$ for $M>1\implies R=0.$ $\Delta R=1$ if $N\geq M$, else $0$
mean Inner: $T_{j}=1$ only if all pick $j$, else 0. Summation $\sum_{j}T_{j}$ is number of tasks chosen by all agents. Usually 0 or 1. Outer: Average across $j$. $R=\frac{1}{M}\sum_{j}T_{j}.$ $\implies R=1/M,$ $\Delta R=0.$ Inner: $T_{j}=\tfrac{|\mathcal{I}_{j}|}{N}$. Outer: Average across tasks: $R=\frac{1}{M}\sum_{j=1}^{M}\frac{|\mathcal{I}_{j}|}{N}=\frac{1}{M}.$ No matter how agents are distributed, $\sum_{j=1}^{M}|\mathcal{I}_{j}|=N$. Hence $R_{\mathrm{het}}=R_{\mathrm{hom}}=\tfrac{1}{M},$ $\Delta R=0.$ Inner: $T_{j}=1$ if chosen by at least 1 agent, else 0. Outer: Average across $j$: $\tfrac{1}{M}\sum_{j}T_{j}.$ This is $\tfrac{1}{M}\cdot\bigl{(}\#\text{ of tasks chosen}\bigr{)}.$ Heterogeneous can pick up to $\min(M,N)$ tasks, so $R=\frac{\min(M,N)}{M}.$ Homogeneous covers exactly 1 task $\implies R=1/M.$ $\Delta R=\frac{\min(M,N)-1}{M}.$
@@
Appendix H Parametrized Families of Aggregators
@@
Power-Sum $f_{t}(\mathbf{x})~{}=~{}\sum_{i=1}^{N}\bigl{(}x_{i}\bigr{)}^{t},\quad x_{i}\geq 0,\quad t>0$ Strictly Schur-convex for $t>1$. Strictly Schur-concave for $10<t<$ 1. At $t=1$, it is linear (both Schur-convex and Schur-concave). Undefined at $t\leq 0$ if any $x_{i}=0$, though one can extend with limits.
Power-Mean $M_{t}(\mathbf{x})~{}=~{}\biggl{(}\frac{1}{N}\sum_{i=1}^{N}(x_{i})^{t}\biggr{)}^{\!\!1/t},\quad x_{i}\geq 0,\;t\neq 0$ Strictly Schur-convex for $t>1$. Strictly Schur-concave for $10<t<$ 1. Reduces to arithmetic mean at $t=1$. As $t\to\infty$, converges to $\max_{i}x_{i}$; as $t\to-\infty$, converges to $\min_{i}x_{i}$.
@@
Appendix I Additional Results in the Multi-Agent Multi-Task Matrix Game
@@ I.3 $N=M=4$
$U$ Min $0.0\pm 0$ 0.0 $0.25\pm 0.0$ $1.0\pm 0.0$
Mean $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.75\pm 0.0$
Max $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
Appendix J Multi-Goal-Capture
In Fig. 7 we juxtapose two representative $N=M=2$ roll-outs of the Multi-Goal-Capture environment for homogeneous teams (top row) and heterogeneous teams (bottom row) when $U~{}=~{}\min,\,T=\max$. Consistent with the discussion in Sec. 5, homogeneous agents steer to the geometric midpoint between the two goals, producing almost overlapping paths–this is suboptimal, as they cannot cover both goals. On the other hand, heterogeneous agents exaggerate their differences–taking sharply diverging trajectories and ensuring one goal each.
@@
Appendix K Parametrized Dec-POMDP

### v2

@@ Appendix A Computational Resources Used
Simply reproucing our results using the available code will take considerably less compute hours (around a day).
@@ Appendix B Code and Data Availability
We attach the code in the supplementary materials. The code contains instructions on how to reproduce the experiments in the paper and dedicated YAML files containing the hyperparameters for each experiment presented. The YAML files are structured according to the HYDRA (Yadan, 2019) framework which allows smooth reproduction as well as systematic and standardized configuration. We further attach all scripts to reproduce the plots in the paper from the experiment results.
@@ Appendix C Implementation Details
For all experiments, we use the MAPPO MARL algorithm (Yu et al., 2022). Environments are implemented in the multi-agent environment simulator VMAS (Bettini et al., 2022), and trained using TorchRL (Bou et al., 2023). Both the actor and critic are two-layer MLPs with 256 neurons per layer and Tanh activation. Further details, such as hyperparameter choices, are available in the attached code and YAML configuration files.
Experiments were run on a single Nvidia L40 GPU. In the HetGPS experiments (Sec. 4), a standard MARL training iteration (60,000 frames) takes approx. 15s; including the environment backpropagation increases this to approximately 20s.
Appendix D Use of LLMs
We used LLMs (ChatGPT 4o and Gemini 2.5 Pro) to improve some parts of the writing, e.g., make wording suggestions. We verified and take responsibility for all LLM-related outputs in this work.
Appendix E Colonel Blotto & Level-Based Foraging
@@
E.1 Team Colonel Blotto (fixed adversary)
The Colonel Blotto game is a well-known allocation game studied in both game theory and MARL (Roberson, 2006; Noel, 2022). It is used to model election strategies and other resource-based competitions. In the team variant with fixed adversary, $N$ friendly colonels (agents) each distribute a (fixed and equal) budget of troops $r_{ij}\!\geq 0,\sum_{j=1}^{M}r_{ij}=1$ across $M$ battlefields $j\in\{1,\dots,M\}$ (our tasks). A fixed adversary selects a stochastic opposing allocation strategy, i.e. a distribution $\pi_{\mathrm{adv}}$ over vectors $a=(a_{1},\dots,a_{M})$ which is fixed throughout training and evaluation. Let $s_{j}=\sum_{i=1}^{N}r_{ij}$ denote the team force committed to battlefield $j$. Our agents win against the adversary if the troops they allocate to a given field surpass the troops allocated by the adversary. The expected value secured on battlefield $j$ is therefore
@@ E.1 Team Colonel Blotto (fixed adversary)
E.2 Level-Based Foraging
The well-known level-based foraging (LBF) benchmark, based on the knapsack problem (Garey et al., 1990), is a deceptively challenging, embodied MARL environment, where $N$ agents are placed on a grid with $M$ food items, and are tasked with collecting them. Each item $j$ has an integer level $L_{j}$ that must be met or exceeded by the combined skills of the agents standing on that cell before it can be collected (Papoudakis et al., 2021). Let agent $i$’s skill be $e_{i}$. At a given step the binary variable
@@ E.2 Level-Based Foraging
Appendix F Sum-Form Aggregators
@@
Definition F.1 (Sum-Form Aggregator).
@@
Lemma F.1 (Schur Properties of Sum-Form Aggregators (Peajcariaac & Tong, 1992)).
@@
Corollary F.1 (Convex-Concave Positive Heterogeneity Gain).
@@
First, by Lemma F.1, since $g$ is strictly convex, the (identical) task-level aggregators $T_{j}(\mathbf{x})=\sum_{i=1}^{N}g(x_{i})$ are strictly Schur-convex, satisfying condition (i) of Theorem 3.1.
@@
Hence, the conditions of Thm. 3.1 apply. To establish $\Delta R>0$, it remains to verify that the optimal allocation is non-trivial: it distributes effort across at least two tasks. In any admissible homogeneous solution, each of the $N$ agents chooses the same effort‐distribution $(c_{1},\dots,c_{M})$ on tasks, with $\sum_{j}c_{j}=1$. Then task $j$’s reward is $T_{j}=N\,g(c_{j})$, so $R(\mathbf{A})\;=\;\sum_{j=1}^{M}h\!\bigl(N\,g(c_{j})\bigr).$ The trivial, all-agent single‐task allocation uses $(c_{j}=1,c_{k\neq j}=0)$. Its reward is therefore $R_{\text{corner}}=h\!\bigl(N\,g(1)\bigr)+\sum_{k\neq j}h\bigl(N\,g(0)\bigr)=h\!\bigl(N\,g(1)\bigr)$ since $g(0)=0$ and $h(0)=0$.
@@
Appendix G Formal Analysis
G.1 Proof of Thm. 3.1
@@ Appendix G Formal Analysis
Then each column $j$ of $\mathbf{A}_{\mathrm{hom}}$ is the uniform vector $\mathbf{u}_{j}\;=\;(c_{j},c_{j},\dots,c_{j})^{\top}\;\in\mathbb{R}^{N}.$ Hence the task‐level reward is $T_{j}(\mathbf{u}_{j})$, and the overall reward is
@@
The total effort allocated to a task $j$ can be expressed as $\lfloor Nc_{j}\rfloor+f_{k}$, where $0\leq f_{j}<1$. First, we assign $\lfloor Nc_{j}\rfloor$ agents to allocate effort $1$ to task $j$, for every task $j$. These agents are all distinct. This leaves us with $\sum_{j}f_{j}~=~N~-~\sum_{j}\lfloor Nc_{j}\rfloor$ agents that have not allocated any effort yet. Let $i$ be the first of those agents. We have agent $i$ allocate $f_{1}$ effort to task $1$, $f_{2}$ effort to task $2$, and so on, until we arrive at a task $k$ such that $f_{1}+\ldots+f_{k}=1+s$, for some $s>0$. We have $i$ allocate $f_{k}-s$ to this task $k$. Then, we move to agent $i+1$, and allocate the remaining fractional efforts in the same manner (and in particular, allocating $s$ effort to task $k$), until agent $i+1$ overflows. Then we move to agent $i+2$, and so on. This ensures that we have allocated $N$ effort in total across the agents, and that every agent’s effort allocation sums exactly to $1$, so is feasible.
@@
which is clearly homogeneous (each row is the same), and respects each column sum $s_{j}$. Since $\sum_{j}s_{j}=N$, each row sums to $1$, hence the allocation is feasible. By Schur‐concavity of $T_{j}$, for each column $j$ we have
@@
unless $\mathbf{a}_{j}$ is $\mathbf{u}_{j}$. In other words, any deviation from the uniform vector with the same sum $\sum_{i=1}^{N}a_{ji}=s_{j}$ will not increase $T_{j}(\mathbf{a}_{j})$ under Schur‐concavity. Hence for each column $j$ of $\mathbf{A}$, $T_{j}(\mathbf{a}_{j})\;\leq\;T_{j}(\mathbf{u}_{j}),$. Since $U$ is non-decreasing in each coordinate,
@@
always sum to $C$. By strict Schur-convexity, the maximum value of $U$ over such vectors is attained precisely at an extreme point of the $C$-simplex, i.e. at some permutation of $(C,0,\ldots,0)$. Hence, we seek to find an allocation of efforts, $\mathbf{A}_{\mathrm{corner}}$, that causes $\textbf{T}(\textbf{A})$ to equal this vector.
@@
$\boxed{\Delta R(t,\tau;N)\geq\begin{cases}\sigma(t,N)-\dfrac{1}{N},&t>0,\;\;\tau\leq 0,\\[10.0pt] \max\!\bigl\{\sigma(t,N)-\sigma(\tau,N),\,0\bigr\},&t>0,\;\;\tau\geq 0.\end{cases}}$
@@
Appendix H Deriving the $\{\min,\textnormal{mean},\max\}$ heterogeneity gains in the Fig. 2 table
@@
$U=$ mean Inner: $T_{j}=\min_{i}r_{ij}=1/M.$ Outer: simple avg $\tfrac{1}{M}\sum_{j}T_{j}.$ Since $\sum_{j}T_{j}=M\cdot(1/M)=1\implies R=1/M.$ Both $R_{het}$, $R_{hom}$ same $\implies\Delta R=0.$ Gap: 0. Inner: $T_{j}=\tfrac{1}{N}\sum_{i}r_{ij}.$ Then $\sum_{j}T_{j}=1.$ Outer: avg $\tfrac{1}{M}\sum_{j}T_{j}.$ Hence $R=\tfrac{1}{M}\cdot 1=\tfrac{1}{M}.$ Same for $R_{het}$, $R_{hom}$. Gap: 0. Inner: $T_{j}=\max_{i}r_{ij}.$ Outer: avg $\tfrac{1}{M}\sum_{j}T_{j}.$ $R_{het}$: sum $M\implies R=1.$ $R_{hom}$: sum $1\implies R=1/M.$ Gap: $1-\tfrac{1}{M}=\tfrac{M-1}{M}.$
@@
[…]
mean Inner: $T_{j}=1$ only if all pick $j$, else 0. Summation $\sum_{j}T_{j}$ is number of tasks chosen by all agents. Usually 0 or 1. Outer: Average across $j$. $R=\frac{1}{M}\sum_{j}T_{j}.$ $\implies R=1/M,$ $\Delta R=0.$ Inner: $T_{j}=\tfrac{|\mathcal{I}_{j}|}{N}$. Outer: Average across tasks: $R=\frac{1}{M}\sum_{j=1}^{M}\frac{|\mathcal{I}_{j}|}{N}=\frac{1}{M}.$ No matter how agents are distributed, $\sum_{j=1}^{M}|\mathcal{I}_{j}|=N$. Hence $R_{\mathrm{het}}=R_{\mathrm{hom}}=\tfrac{1}{M},$ $\Delta R=0.$ Inner: $T_{j}=1$ if chosen by at least 1 agent, else 0. Outer: Average across $j$: $\tfrac{1}{M}\sum_{j}T_{j}.$ This is $\tfrac{1}{M}\cdot\bigl(\#\text{ of tasks chosen}\bigr).$ - Heterogeneous can pick up to $\min(M,N)$ tasks, so $R=\frac{\min(M,N)}{M}.$ - Homogeneous covers exactly 1 task $\implies R=1/M.$ $\Delta R=\frac{\min(M,N)-1}{M}.$
@@
Appendix I Parametrized Families of Aggregators
@@
Power-Sum $f_{t}(\mathbf{x})~=~\sum_{i=1}^{N}\bigl(x_{i}\bigr)^{t},\quad x_{i}\geq 0,\quad t>0$ Strictly Schur-convex for $t>1$. Strictly Schur-concave for $0<t<1$. At $t=1$, it is linear (both Schur-convex and Schur-concave). Undefined at $t\leq 0$ if any $x_{i}=0$, though one can extend with limits.
Power-Mean $M_{t}(\mathbf{x})~=~\biggl(\frac{1}{N}\sum_{i=1}^{N}(x_{i})^{t}\biggr)^{\!\!1/t},\quad x_{i}\geq 0,\;t\neq 0$ Strictly Schur-convex for $t>1$. Strictly Schur-concave for $0<t<1$. Reduces to arithmetic mean at $t=1$. As $t\to\infty$, converges to $\max_{i}x_{i}$; as $t\to$-$\infty$, converges to $\min_{i}x_{i}$.
@@
Appendix J Additional Results in the Multi-Agent Multi-Task Matrix Game
@@ 3 $r_{31}$ $r_{32}$ $r_{33}$
$U$
Min $0.0\pm 0.0$ $\bm{0.25}\pm 0.0$ $\bm{1.0}\pm 0.0$
Mean $\pm 0.0 0.0±0.0$$\pm 0.0$ $\bm{0.75}\pm 0.0$
Max $\pm 0.0 0.0±0.0$$\pm 0.0$ $0.0\pm 0.0$
J.4 $N=M=8$
To test scalability, we further report results for the discrete matrix game with $N=M=8$ in Tab. 9. As discussed in Sec. 5, due to the larger agent scale and action dimensionality (8 agents and 8 tasks), for some reward structures the empirical heterogeneity gain is negative, since the neurally heterogeneous agents we train require more time to discover the optimal policy, hence lag behind their homogeneous counterparts under a fixed training budget. Increasing the number of training steps steers the negative $\Delta R$ values to 0. This nuance aside, what is important is that the results still follow our theory exactly in terms of which reward structures yield a positive heterogeneity gain. For such reward structures, our theory also precisely predicts the numerical value of $\Delta R$ (Fig. 2) .
Table 9: Heterogeneity gain $\Delta R\in\mathbb{R}_{0\leqslant x\leqslant 1}$ of the discrete matrix game with $N=M=8$. We report mean and standard deviation after 12 million training frames over 8 different random seeds.
$T$
Min Mean Max
$U$
Min $0.0\pm 0.0$ $\bm{0.125}\pm 0.0$ $\bm{1.0}\pm 0.0$
Mean $94\pm 0.068 0.0±0.0$$\pm 0.0$ $\bm{0.875}\pm 0.0$
Max $75\pm 0.5 0.0$$\pm 0.0 0.0$$0\pm 0.0$
Appendix K Multi-goal-capture
In Multi-goal-capture, agents need to navigate to goals. Each agent observes the relative position to the goals, and agent actions are continuous 2D forces that determine their direction of motion. The entries $r^{t}_{ij}$ of matrix $\mathbf{A}^{t}$ at time $t$ represent the local reward of agent $i$ towards goal $j$, computed as $r^{t}_{ij}=\left(1-d^{t}_{ij}/\sum_{j=1}^{M}d^{t}_{ij}\right)/(M-1),$ where $d^{t}_{ij}$ is the distance between agent $i$ and goal $j$. This makes it so that $\sum_{j=1}^{M}r^{t}_{ij}=1$ and $r^{t}_{ij}\geqslant 0$. At each step, the agents receive the global reward $R(\mathbf{A}^{t})$, with aggregators $U,T\in\{\min,\mathrm{mean},\max\}$.
Our results, shown in Fig. 4(a), show that our curvature theory reliably predicts when there is a heterogeneity gain (Fig. 2): it is positive only for the concave–convex pairs $U=\min,\,T=\max$ and $U=\mathrm{mean},\,T=\max$. The heterogeneity gain is smaller in the latter case because learning dynamics matter: with $U=\min,\,T=\max$ the best homogeneous policy is unique (every agent must steer to the midpoint between the two goals) so homogeneous learners seldom find it, leaving room for heterogeneous policies to excel (see App. K). By contrast, $U=\mathrm{mean},\,T=\max$ admits a continuum of good homogeneous policies, which homogeneous teams execute more easily. For $U=\max,\,T=\min$ and $U=\max,\,T=\mathrm{mean}$, the theoretical $\Delta R$ is 0, yet the empirical heterogeneity gap is negative (4(a)). This occurs because the reward peaks only when all agents coordinate on the same goal. Neurally heterogeneous teams learn this uniform behavior slower than homogeneous teams, so they underperform within the fixed training budget. Additional training would close this gap to $\Delta R=0$.
In Fig. 7 we juxtapose two representative $N=M=2$ roll-outs of the Multi-goal-capture environment for homogeneous teams (top row) and heterogeneous teams (bottom row) when $U~=~\min,\,T=\max$. Consistent with the discussion in Sec. 5, homogeneous agents steer to the geometric midpoint between the two goals, producing almost overlapping paths–this is suboptimal, as they cannot cover both goals. On the other hand, heterogeneous agents exaggerate their differences, taking sharply diverging trajectories and ensuring one goal each.
@@
Appendix L 2v2 Tag Experiments
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Comparison of homogeneous (top row) and heterogeneous (bottom row) 2v2 tag policies for chaser agents, trained with the reward structure $U=\min,T=\max$ across different initializations. Every column shows the trajectory of the homogeneous (top) and heterogeneous (bottom) policies. (Note that trajectories here are smoothened; agents don’t go over obstacles in actual execution). The heterogeneous policies prioritize capturing both agents, whereas the homogeneous policies focus on just one. In the $U=\min,T=\max$ setting, this gives heterogeneous agents greater reward, hence $\Delta R>0$. Please find more visualizations on our website.
The goal of our tag experiment is to showcase that our theoretical results, which predict the value of $\Delta R$ based on the curvature of the aggregators, hold for discrete, sparse rewards. Specifically, our results for discrete efforts in Fig. 2 predict that only $(U,T)=(\min,\max),(\min,\text{mean}),(\text{mean},\max)$ will have positive heterogeneity gain, with $(\min,\max)$ maximizing the gain. We show in 4(b) that this holds in 2v2 tag, despite the fact that this is a challenging, embodied, long-horizon, whereas our formal results are for instantaneous allocation games [^4] The gain for $(\text{mean},\max)$ is small compared to the other two aggregator combinations, but still positive at $\Delta R\approx 0.37$. This is also significantly higher than aggregator combinations for which we predict $\Delta R$ is not positive, the largest of which attained $\Delta R<0.01$.. Note that this is a highly interpretable result: a $(\min,\max)$ means that agents are only awarded when both escapers are caught, incentivizing heterogeneous strategies where chaser agents split their behaviour so that the chasing efforts $r^{t}_{ij}$ are equally distributed between both escapers. Fig. 8 visualizes the trajectories learned by agents trained under this $(\min,\max)$ reward structure, showing distinct emergent pursuit strategies emerging depending on whether the agents are neurally heterogeneous or neurally homogeneous.
Appendix M Football Experiments
Table 10: Football heterogeneity gains across different reward formulations. Results obtained after 500 training iterations of 240k frames each (6 seeds). Opponent speed annealed from 0% to 100%.
Reward $\Delta R$ Theory $\Delta R>0$? Reward meaning
$U=\min,\;T=\max$ $\bm{1.76}\pm 0.72$ Yes One agent should attend the ball, the other the opponent; reward capped by the less-covered task.
$U=\text{mean},\;T=\max$ $\bm{1.18}\pm 0.11$ Yes Similar to $(\min,\max)$, but reward is dictated by average task performance.
$U=\text{mean},\;T=\text{mean}$ $0.009\pm 0.075$ No Agents should attend both the opponent and the ball.
$U=\min,\;T=\min$ $0.08\pm 0.73$ No At least one agent should attend at least the opponent or the ball.
In some environments, the reward structure might not entirely follow the double-generalized-aggregator structure we study in this work, but at least some part of the reward function might obey this structure. In our study of the VMAS football scenario (Bettini et al., 2022), we ask what happens when this is the case.
Football is a complex, embodied, long-horizon scenario that requires the agents to learn low-level dribbling skills as well as high-level strategy purely from a shared cooperative reward. The VMAS scenario uses reward shaping to enable agents to learn such behaviors. We add a reward structure $U(T(r_{11}^{t},r_{21}^{t}),T(r_{12}^{t},r_{22}^{t}))$ on top of this and ask how this affects heterogeneity.
In our experimental scenario, two learning agents spawn at midfield. A ball is located between them and the goal to the right; a heuristic defender spawns to their left and chases the ball. Agents receive a global reward that increases when the ball moves toward the goal and the defender stays away from it. Additionally, we reuse the reward structure from our Multi‑Goal‑Capture to define rewards for two tasks: tackling the ball, and tackling the opponent.
The effort at time $t$ is:
$r_{ij}^{t}=(1-\frac{d_{ij}^{t}}{\sum_{j}d_{ij}^{t}})/d_{ij}^{t},$
where $d_{ij}$ is distance of agent $i$ to ball or opponent). The global reward given to all agents is then computed as:
$R^{t}=U(T(r_{11}^{t},r_{21}^{t}),T(r_{12}^{t},r_{22}^{t}))+\beta[(d_{ball,goal}^{t-1}-d_{ball,goal}^{t})-(d_{opp,ball}^{t-1}-d_{opp,ball}^{t})],$
where $\beta$ weighs the global football reward.
Since this reward structure does not follow our theory entirely, we ask whether, when $U,T$ are, respectively, strictly Schur-concave and strictly Schur-convex, we should expect $\Delta R>0$ as in our other scenarios. We test this for $U=\min,T=\max$ and $U=\text{mean},T=\max$. To control for the possibility that football is heterogeneous “by default”, we also test the aggregator combinations $U=\text{mean},T=\text{mean}$ and $U=\min,T=\min$ as controls.
We report heterogeneity gains after training homogeneous and heterogeneous policies in Tab. 10. This shows that our curvature test predicts the heterogeneity gain of different reward structures, despite only being a component in the overall reward structure. This insight is important, as it indicates our theoretical insights (the curvature test) may extend beyond environments that strictly follow our task allocation setting.
The resulting policies are reported in Fig. 9 with videos here.
Figure 9: Left: Results of training a heterogeneous policy on VMAS Football where agents are trained with $U=\min,T=\max$ aggregators. Our learning agents are drawn in blue; the heuristic opponent in red; and the ball in black. The learning agents split their efforts, tackling both the ball and the opponent. Right: Results of training a homogeneous policy. Agents are unable to split their efforts, so either they both tackle the ball, or both tackle the opponent. This results in lower reward, hence $\Delta R>0$. These policies are visualized on our website.
Appendix N Observability-Heterogeneity Trade-Off
Figure 10: Gain w.r.t. observability when $U=\min,T=\max$.
Figure 11: Heterogeneity gain for Multi-goal-capture throughout training when the agents’ observation range is gradually increased from 0 to $0.35$ over 4 random seeds (4 random seeds suffice as this phenomenon is established in the literature (Bettini et al., 2023), and we only wish to show its emergence in the context of our work.)
In this Appendix, we crystallize the relationship between environment observability and empirical heterogeneity gains. It is well known that neurally homogeneous agents (i.e., sharing the same parameters) can achieve behavioral heterogeneity by conditioning their actions on diverse input contexts (behavioral typing). This can be achieved by naively appending the agent index to its observation (Gupta et al., 2017a) or by providing relevant observations that allows the agents to infer their role (Bettini et al., 2023). Behavioral typing is impossible in matrix games, as these games are observationless. However, it is possible in more complex games, such as our Multi-goal-capture scenario. We augment agents in the positive gain scenario ($U=\min,T=\max$) with a range sensor, providing proximity readings for other agents within a radius. In Fig. 11, we show that the heterogeneity gain decreases as the agent visibility increases (higher sensing radius). This is because, with a higher range, homogeneous agents can sense each other and coordinate to pursue different goals. This result highlights the tight interdependence between the heterogeneity gain and agents’ observations.
Appendix O Parametrized Dec-POMDP